OpenStudy (anonymous):

Write log(8) 3 as a logarithm of base 5.

OpenStudy (anonymous):

I said \[\frac{ \log _{8}5}{ \log _{3}5}\]

OpenStudy (imstuck):

there is a conversion formula. Let me type it for you.

OpenStudy (anonymous):

Thank you!

OpenStudy (imstuck):

the formula is this:\[\log _{b}M=\frac{ \log _{a}M }{ \log _{a}b }\]

OpenStudy (anonymous):

Ohhhh.. Thank You! I get it.

OpenStudy (imstuck):

your answer would be

OpenStudy (imstuck):

\[\log _{8}3=\frac{ \log _{5} 3}{ \log _{5}8 }\]you could use your calculator now to find the exact value of that, if you want/need to.